The results of an independent-measures study produce a t statistic with df = 36. How many individual participated in the entire study?
Answer #1

The degree of freedom of independent-measures study with sample size n1 and n2 is

df = n1+n2-2

Here we need to find n=n1+n2 so

36 = n -2

n = 36+2 = 38

That is total individual participated in the entire study is 38.